A bottling plant fills 2,400 bottles every two hours. The lead time is 50 minutes and a container accommodates 80 bottles. The safety stock is 30 percent of expected demand. How many kanban cards are needed? (Round up your answer to the next whole number.)
Answer #1

Number of kanbans = DL(1 + S)/C

D = Average number of units demanded over some time period

L = Lead time to replenish an order (expressed in the same units as demand)

S = Safety stock expressed as a percentage of demand during the lead time

C = Container size

D = If the average number of units demanded is 2400 and the time period is 2 hours, then that's the same as 1200 in an hour, 1200 in 60 minutes, 20 in one minute.

L = 50

S = 0.3

C = 80

Number of kanbans = 20 * 50 (1 + 0.3) / 80 = 16.25

Do give a positive rating if you think this answer helped